### v3

# UDA4Inst: Unsupervised Domain Adaptation for Instance Segmentation

Abstract
Unsupervised Domain Adaptation (UDA) aims to transfer knowledge learned from a labeled source domain to an unlabeled target domain. While UDA methods for synthetic to real-world domains (synth-to-real) show remarkable performance in tasks such as semantic segmentation and object detection, very few were proposed for instance segmentation in the field of vision-based autonomous driving, and the existing ones are based on a suboptimal baseline, which severely limits the performance. In this paper, we introduce UDA4Inst, a strong baseline of synth-to-real UDA for instance segmentation. UDA4Inst adopts cross-domain bidirectional data mixing at the instance level to effectively utilize data from both source and target domains. Rare-class balancing and category module training are also employed to further improve the performance. It is worth noting that we are the first to demonstrate results on two new synth-to-real instance segmentation benchmarks, with 39.0 mAP on UrbanSyn Cityscapes and 35.7 mAP on Synscapes Cityscapes. Our method outperforms the source-only Mask2Former model by +7 mAP and +7.6 mAP, respectively. On SYNTHIA Cityscapes, our method improves the source-only Mask2Former by +6.7 mAP, achieving state-of-the-art results.111Our code will be released soon.
Keywords:
Synth-to-Real Unsupervised Domain Adaptation Instance Segmentation Vision-based Autonomous Driving.1 Introduction
Instance segmentation provides a detailed and fine-grained understanding of the objects in an image, which is critical for a variety of real-world applications, such as multi-object tracking [24, 41, 2], pose estimation [6, 1], and motion prediction [51, 29]. Specifically, in autonomous driving, instance segmentation is useful as it helps to locate safety-critical objects (such as pedestrians or riders) with accurate boundaries, enabling more advanced scene analysis and decision-making in many scenarios [9, 49, 2, 52].
However, in practice, training instance segmentation models requires access to a large amount of densely labeled data. Each instance needs to be accurately delineated with pixel-level masks, which is laborious and highly time-consuming. In addition, to ensure proper generalization of the model, the data should contain a wide range of object appearances, poses, scales, and backgrounds. Therefore, the label acquisition and object variability needed for instance segmentation encounter a data-hungry bottleneck. This challenge demands a more elaborate approach to enhance the efficiency of the labeling of instance segmentation data.
As an alternative, synthetic data are increasingly being used to train deep models as they enable large-scale training data, which is especially relevant for vision tasks [38]. Synthetic data offers labels annotated directly by simulators for free [10, 37]. Crucially, it allows for the incorporation of diversity into the training dataset, thereby enhancing the model’s generalization capability.
However, when training models with synthetic data and testing them in the real-world domain, the results are often unsatisfactory due to the inevitable gap between the two domains. UDA methods are essential for mitigating this inherent gap between the synthetic source and the real-world target domains while avoiding human labeling. In recent years, many studies have explored the integration of synthetic data and UDA methods to address a wide range of computer vision tasks including image classification [11, 33, 40, 35], object detection [46, 13, 15, 39, 54, 36], and semantic segmentation [3, 19, 20, 23, 16, 17, 32]. Comparatively, there has been relatively little research on UDA for instance segmentation, reflected in only three works [55, 22, 44] in the last five years. Among them, although [22, 44] do provide results on instance segmentation, they were models designed for the panoptic segmentation task, and [55] was based on a suboptimal Mask R-CNN framework with adversarial learning.
In this paper, we focus on designing a framework that provides a strong baseline on UDA for Instance Segmentation (UDA4Inst) following an online self-training pattern. Unlike previous models [11, 20, 24] that perform typical unidirectional cross-domain data mixing, i.e., either source-to-target (S2T) or target-to-source (T2S), UDA4Inst utilizes an instance-based bidirectional cross-domain data mixing approach tailored for instance segmentation. For the T2S phase, we need to take pseudo labels for real-world instances and accommodate them into synthetic images. When multiple instances of the same class overlap, in semantic segmentation, these overlapping instances are treated as a single blob. In contrast, instance segmentation methods need to find clear boundaries in the overlapping region of objects to separate them. Note that this is especially more difficult than semantic segmentation. Additionally, to mitigate the problem of instance imbalance in existing datasets, UDA4Inst incorporates a handy online algorithm to balance rare-class instances during data mixing between domains. In this way, we observe a clear improvement in the segmentation performance of rare-class instances. As the third methodological contribution, we explore category module training, where multiple classes are categorized into several modules for separate training, then we concatenate outputs of these modules to obtain a final prediction. We demonstrate the effectiveness of UDA4Inst on three UDA instance segmentation benchmarks. To the best of our knowledge, we are the first to perform UDA instance segmentation using the synthetic driving scene datasets UrbanSyn [14] and Synscapes [53], achieving 39.0 mAP and 35.7 mAP on Cityscapes [9]. Additionally, we have achieved state-of-the-art results on SYNTHIA [42] Cityscapes. Compared to the source-only method, our approach shows improvements of +7 mAP, +7.6 mAP, and +6.7 mAP on these three benchmarks.
2 Related Work
2.1 Instance Segmentation
Early mainstream deep learning models based on convolutional neural networks (CNNs) for instance segmentation, such as Mask R-CNN [18] and YOLACT [4], have achieved remarkable successes. In recent years, with the adoption of the self-attention mechanism in natural language processing and computer vision, transformer-based models have emerged and surpassed CNN-based architectures in many tasks such as object detection [28], semantic segmentation [8, 7], and panoptic segmentation. [44, 7]. Among them, Mask2Former [7] has attracted great attention as a universal transformer-based architecture that obtains top results in three major image segmentation tasks (panoptic, instance, and semantic) on several benchmarks. It introduces masked attention, which extracts localized features by constraining cross-attention within predicted mask regions. In this work, we leverage this advanced universal segmentation architecture to prompt the state-of-the-art results in instance segmentation with UDA.
2.2 Unsupervised Domain Adaptation
Typically, UDA methods occupy three groups: domain discrepancy alignment, adversarial learning, and self-training. In domain discrepancy alignment, a proper divergence measure, such as maximum mean discrepancy [43], correlation alignment [47], or Wasserstein distance [31] is chosen to minimize domain discrepancy in a latent feature space. Adversarial learning approaches [45, 50] aim to achieve domain invariance in input space, feature extraction, or output space. To reduce domain discrepancy, self-training methods [19, 20, 21] train a model using a combination of labeled and unlabeled data. These methods generate high-confidence pseudo-labels from unlabeled data and iteratively incorporate them to expand the labeled data during training, gradually improving the model’s performance. HRDA [20] enhances domain adaptation performance through the integration of hierarchical attention and multi-resolution representations. MIC [21] leverages masked image consistency to enhance context understanding for generalization. Our proposed solution follows the pattern of self-training since it is the mainstream approach in state-of-the-art UDA methods, and is most suitable for our bidirectional mix strategy.
Current UDA works [19, 20, 23, 16, 17, 32] mainly focus in the semantic segmentation task that is getting really close to the supervised learning methods trained on human labeling datasets. On the other hand, UDA panoptic segmentation is yet to be thoroughly explored, where few works [22, 44] can simultaneously achieve domain-adaptive instance and semantic segmentation within a single network. The work most related to ours is [55], which proposes a multi-level feature adaptation strategy to improve synth-to-real instance segmentation. To reduce the domain gap, they focus on aligning feature distributions between synthetic and real data at global, local, and subtle levels. This is achieved by using Mask R-CNN and discriminator networks with adversarial learning. In recent years, there has been significant progress in the research fields of non-cross-domain instance segmentation and UDA methods. In this work, we aim to improve the baseline on UDA for instance segmentation with approaches such as cross-domain data mixing along with the advanced transformer-based Mask2Former network [7].
2.3 Cross-domain Data Mixing
A commonly used data augmentation technique is mixing image content. In UDA, this augmentation is usually applied in the form of cross-domain mixing, i.e. the mixed samples consist of data from both source and target domains. Tranheden et al. introduce DACS [48], a unidirectional cross-domain data mixing strategy that cuts out semantic classes in the source domain and pastes them into samples in the target domain. Suhyeon et al. use a similar method by transferring the tail-class content from source to target to solve the class imbalance problem [27]. Gao et al. propose a dual soft-paste strategy that not only considers mixing from the source to target images but also within the source domain itself [12]. A similar idea is suggested by Huo et al., while for the dual mixing [25], they focus on the target-to-target domain. Most recently, Kim et al. propose a S2T and T2S cut-and-paste strategy based on predefined image patches [26]. It is noteworthy that these mixing approaches fall into two categories: 1) Class-wise Mixing. For instance, Cardace et al. [5] only mix semantic classes such as ‘Car’, ‘Person’, and ‘Pole’. This approach fails to delineate boundaries accurately where multiple instances overlap. 2) Random Patches Mixing which involves cut-and-paste based on arbitrary patches, resulting in unrealistic images with mixed foreground instances and background area.
Unlike the prior works in semantic segmentation, we propose a bidirectional domain mixing strategy tailored for instance segmentation, which involves both S2T and T2S mixing. Since our work focuses on instance segmentation, we perform cut-and-paste based on semantic instances rather than classes or random patches. Therefore, instances pasted in the new domain do not include background pixels from the original domain, generating more realistic mixed data.
3 Methodology
3.1 Overall Approach
For the source domain dataset , comprising images and their instance labels , an instance segmentation model can be trained in a supervised manner. However, the target domain dataset only provides unlabeled images . We aim to address the domain gap by leveraging , , and to train a model capable of accurately segmenting instances on .
As shown in Fig. 1, our UDA framework is based on online self-training. The training has two stages. In the first stage, the model with parameters is trained on the source domain in a supervised manner for iterations, with synthetic images and their corresponding ground-truth labels . Our goal is to minimize a supervised segmentation loss
then the network can be used to predict pseudo-labels, , on the target domain. In our case, we use Mask2Former [7] as the base network for instance segmentation. The segmentation loss for Mask2Former is
where binary cross-entropy loss and dice loss are for binary mask segmentation, is for instance class classification. , and are weighting hyper-parameters.
In the second stage, we utilize the cross-domain bidirectional mixed images and labels produced by both source and target domain data to resume the training of the model for iterations. The rare-class synthetic instances defined in the source domain are resampled during data mixing. Since the target images are unlabeled, we utilize the pseudo-labels generated by the exponential moving average (EMA) teacher model . At the beginning of the second stage, the parameters of the EMA teacher are initialized by the weights of . Then, in each training step, we first infer the target images by the EMA teacher to get pseudo-labels . In line with common practice [48, 20, 19, 21], we use confidence thresholds to filter out noisy pseudo-labels. A pseudo-label confidence score is calculated as the product of mask confidence (confidence of whether an instance object exists in a mask) and class confidence (confidence of an instance class). Then, we leverage the ground-truth labels and the pseudo-labels to perform bidirectional instance cut-and-paste, thereby obtaining mixed data . Finally, mixed data are used to train the model with a weighted bidirectional data mixing loss
where and are weights for different directional mixing data. Then the EMA teacher parameters are updated. We present the algorithm process in the supplementary material.
3.1.1 Bidirectional Data Mixing.
In data mixing, the mixed images are generated by cut-and-paste image regions from one image to another. We adopt a bidirectional mixing strategy that fully utilizes instances from both source and target domains to facilitate instance segmentation. As shown in Fig. 2, in each iteration of the mixed data training stage, the mixed data is generated in both S2T and T2S directions. For S2T mixing, the instances in a source domain image are cut and pasted onto a target domain image , resulting in an augmented mixing images . Its corresponding instance label is created according to the image, where the instance labels are added to the generated target pseudo-labels . Similarly, to obtain the T2S mixing data and , we leverage the generated pseudo-labels of the target images to mask out the instances that are desired to be pasted on the source image . The network is trained on mixed images , with mixed labels , .
3.1.2 Category Module Training.
Conventionally, the instance labels have specific semantic meanings such as: ‘Motorcycle’, ‘Bicycle’, ‘Train’ etc. Instances with similar semantics tend to have more common features. For example, ‘Motorcycle - Bicycle’ may contain more common features than ‘Train - Bicycle’, therefore leading to a higher inter-class correlation. Based on this inductive bias, we propose category module training. Instead of training one model to predict all classes, we separate the classes into several category groups. We train one model for each group and combine the outputs of these models to obtain the final result. We assume that introducing this inductive bias may guide the model to better learn and associate the inter-class and intra-class features. To verify the idea, we provide rich ablation study results in Sec. 4.2.4. In this work, we consider the number of category groups as two.
The category module training can greatly improve the accuracy, yet there is a trade-off between model performance and testing efficiency, as it requires two category modules for inference. Considering the testing efficiency, we also provide an alternative that only uses one model during testing, which can achieve comparable performance. In this way, we use the two-category models to generate pseudo-labels during the self-training phase and train a single model which is then used for testing. We name this method as category-free and provide results in the Sec. 4.2.3.
3.1.3 Rare-class Balancing.
Class imbalance is a common issue that harms the performance of deep learning models. Models trained on imbalanced datasets tend to favor classes with more data, leading to poor performance in classes with fewer examples. This problem exists in the three synthetic datasets we used. Therefore, we increase the training frequency of these rare class instances by intentionally resampling them in data mixing. In each category group, for the class that occupies the smallest percentage of instances in the source dataset, we define it as a rare class: ‘Motorcycle’ and ‘Train’ for UrbanSyn (2.47% and 0.37%) and Synscapes (5.91% and 2.55%), ‘Motorcycle’ and ‘Bus’ for SYNTHIA (2.18% and 2.08%).
Specifically, during the process of traversing images for data mixing, for those images containing rare-class instances, we store these images and the corresponding instance labels in a pool with a maximum number of N. These samples are continuously updated with newly emerged rare-class instances and their corresponding images in a FIFO (first-in, first-out) fashion. In turn, when encountering images that do not contain rare classes, we randomly select half of the samples in the pool and mix them on these images. To avoid error accumulation from pseudo-labels, we only apply rare-class balancing to data mixing in the S2T direction. Note that unlike [21], we do not need to pre-process data to gather information about the distribution of instances for each class in the dataset.
| AP | AP50 | |||||||||||||||||||
| Model | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP | Imp. | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP50 | Imp. |
| upper bound | 41.1 | 32.4 | 61.4 | 44.8 | 67.5 | 50.7 | 26.8 | 26.6 | 43.9 | - | 75.4 | 71.3 | 87.2 | 58.7 | 80.3 | 76.1 | 58.1 | 65.1 | 71.5 | - |
| UrbanSyn Cityscapes | ||||||||||||||||||||
| w/o UDA | 25.1 | 19.5 | 39.8 | 37.6 | 51.5 | 46.7 | 19.2 | 16.2 | 32.0 | - | 52.0 | 48.4 | 61.5 | 48.6 | 66.2 | 70.0 | 42.6 | 46.0 | 54.4 | - |
| UDA4Inst | 30.0 | 26.4 | 49.2 | 42.9 | 64.4 | 58.1 | 21.3 | 19.7 | 39.0 | +7.0 | 60.1 | 60.7 | 75.1 | 56.4 | 78.0 | 77.8 | 48.7 | 53.6 | 63.8 | +9.4 |
| Synscapes Cityscapes | ||||||||||||||||||||
| w/o UDA | 27.2 | 19.2 | 39.5 | 24.0 | 46.7 | 33.1 | 18.7 | 16.6 | 28.1 | - | 53.9 | 44.3 | 67.8 | 31.9 | 60.6 | 58.1 | 39.9 | 45.4 | 50.2 | - |
| UDA4Inst | 30.3 | 27.1 | 47.5 | 34.5 | 58.5 | 46.3 | 22.3 | 19.1 | 35.7 | +7.6 | 59.9 | 63.2 | 77.9 | 45.3 | 72.8 | 63.5 | 46.1 | 51.7 | 60.1 | +9.9 |
| SYNTHIA Cityscapes | ||||||||||||||||||||
| Zhang [55] w/o UDA | 7.5 | 0.6 | 19.1 | - | 17.0 | - | 2.5 | 0.7 | 7.9 | - | 25.2 | 4.3 | 39.3 | - | 31.3 | - | 9.6 | 4.0 | 18.9 | - |
| Zhang [55] | 9.4 | 3.8 | 22.2 | - | 23.5 | - | 3.2 | 2.4 | 10.8 | +2.9 | 28.8 | 17.8 | 39.9 | - | 37.5 | - | 12.8 | 11.9 | 24.8 | +5.9 |
| CVRN [22] w/o UDA | - | - | - | - | - | - | - | - | - | - | 30.4 | 9.2 | 32.5 | - | 20.4 | - | 3.1 | 1.3 | 16.2 | - |
| CVRN [22] | - | - | - | - | - | - | - | - | - | - | 34.4 | 25.3 | 38.7 | - | 38.1 | - | 10.1 | 8.7 | 25.9 | +9.7 |
| EDAPS [44] w/o UDA | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 22.0 | - |
| EDAPS [44] | 23.5 | 13.4 | 27.5 | - | 23.6 | - | 5.8 | 0.1 | 15.7 | - | 53.9 | 40.6 | 50.5 | - | 38.8 | - | 21.1 | 0.7 | 34.3 | +12.3 |
| w/o UDA | 26.3 | 15.2 | 38.3 | - | 43.0 | - | 13.6 | 10.9 | 24.6 | - | 53.4 | 40.8 | 68.1 | - | 56.5 | - | 36.4 | 36.1 | 48.6 | - |
| UDA4Inst | 28.1 | 19.6 | 47.9 | - | 58.7 | - | 18.5 | 15.0 | 31.3 | +6.7 | 57.7 | 52.5 | 76.9 | - | 68.6 | - | 42.6 | 43.3 | 56.9 | +8.3 |
4 Experiments
4.1 Implementation Details
we conduct experiments in this work in three synth-to-real domain settings: UrbanSyn Cityscapes, Synscapes Cityscapes, and SYNTHIA Cityscapes.
4.1.1 Training
We adopt Mask2Former [7] with Swin Transformer backbone [32] as our base network . The model is pre-trained on COCO [30] dataset. To minimize the color discrepancy, we adopt the same image pre-processing as [17], aligning the style of source domain images to the target ones by shifting the distribution of pixel values in CIELAB color space. We apply the same data augmentation (flipping, random cropping, etc.) as [7] and set the image crop size to . The maximum number of samples N for the rare-class pool is set to 10, and the pseudo-label confidence threshold is 0.9. We set the numbers of iterations for source domain training to 15k, while the mixed data training is 25k iterations. During validation, we follow the same checkpoint selection protocol used in [21]. For each stage of training, we set a batch size =3 considering optimal memory usage. We use AdamW [34] optimizer with an initial learning rate of 0.0001 and a weight decay of 0.05, as well as the ‘poly’ learning rate schedule with the power of 1. The hyper-parameters of loss , , and are set to 5.0, 5.0, 2.0 respectively. For three synth-to-real settings, we set the weights of bidirectional data mixing loss , to: 0.4, 0.6 for UrbanSyn Cityscapes and SYNTHIA Cityscapes, 0.75, 0.25 for Synscapes Cityscapes. For these experiments, UDA4Inst runs on 3 NVIDIA A40 GPUs, each with 48GB of memory.
4.1.2 Evaluation
We evaluate the UDA performance of 8-class instance segmentation on the validation set of Cityscapes. Following the conventional evaluation criteria of instance segmentation, we take AP (Average Precision) and AP50 (Average Precision at IOU>0.5) and their mean values as our metrics for comparison. All the numbers present in the tables are in %.
(a) AP
(b) AP50
4.2 Experimental Results
4.2.1 UDA performance.
In Table 1, we summarize the results of models tested on the Cityscapes validation set. In three synth-to-real settings, we show results of the UDA4Inst model and the source-only vanilla Mask2Former model [7]. To establish an upper bound, we retrain the Mask2Former model on the Cityscapes training set, i.e., using human-provided labels, reproducing similar results in their paper [7]. Overall, the first observation is that in all settings, UDA4Inst exhibits a significant improvement in both mAP and mAP50 compared to the source-only model. The most evident improvement is reflected in the setting of Synscapes Cityscapes, where the mAP and mAP50 increased by +7.6 and +9.9 points respectively. Our model achieves the highest mAP (39.0) in UrbanSyn Cityscapes, 35.7 mAP in Synscapes Cityscapes, and 31.3 mAP in SYNTHIA Cityscapes. Note that we are the first to conduct experiments in the first two settings for instance segmentation. We observe that models show better adaptation performance in UrbanSyn Cityscapes. Notably, in this setting, UDA4Inst achieves 58.1 AP and 77.8 AP50 on the class of ‘Train’, which surpasses the upper bound model with gains of +7.4 and +1.7. UDA4Inst shows very close AP and AP50 results to the upper bound for ‘Truck’ and ‘Bus’ classes. We hypothesize that these results reflect a smaller domain gap between UrbanSyn and Cityscapes. In Fig. 4, we visualize the predictions and the error maps of the source-only model and UDA4Inst for this setting. We provide more visualizations in the supplementary material. In particular, UDA4Inst demonstrates consistent enhancement over the source-only model across all the classes in all synth-to-real settings. We provide bar charts of class-wise AP and AP50 results in Fig. 3 to see the gains of UDA4Inst over the source-only one.
| Mixing Strategy | mAP |
| w/o mixing | 31.95 |
| S2T mixing | 36.07 |
| T2S mixing | 35.00 |
| bidirectional mixing | 37.18 |
| mAP | ||
| - | - | 34.80 |
| 0.75 | 0.25 | 38.76 |
| 0.5 | 0.5 | 38.09 |
| 0.4 | 0.6 | 38.99 |
As a reference, at the bottom of the table, we list results from Zhang [55], CVRN [22], and EDAPS [44] that were proposed in recent years to address UDA for instance segmentation in SYNTHIA Cityscapes setting. Considering that the listed models used different base networks (i.e. either Mask R-CNN, DAFormer, or Mask2Former), we show the relative improvements in these models after implementing UDA techniques in Table 1. For the same synth-to-real setting, the UDA method proposed by Zhang improves the original model (w/o UDA) by +2.9 mAP and +5.9 mAP50. In contrast, our method achieves improvements of +6.7 mAP and +8.3 mAP50, respectively, demonstrating the superior effectiveness of the UDA approach in our framework. CVRN increases mAP50 from 16.2 to 25.9 (+9.7), and EDAPS increases it from 22.0 to 34.3 (+12.3). Our method, on the other hand, boosts it from 48.6 to 56.9 (+8.3). Even though our mAP50 gain is slightly lower, we work in a significantly higher performance regime, where improvement is harder to obtain.
4.2.2 Bidirectional Data Mixing.
To illustrate the advantages of the Bidirectional Data Mixing (BDM), we conduct experiments on different data mixing strategies without applying Rare-class Balancing (RcB) and Category Module Training (CMT). The results in the UrbanSyn Cityscapes setting are shown in Table 4.2.1. The mAP results reveal that both S2T and T2S mixing contribute to better performance. T2S is slightly inferior to S2T, which we believe is because the noise in the target pseudo-labels can inevitably cause interference in data mixing. Combining S2T and T2S mixing achieves the best result of 37.18 mAP, +5.23 higher than the plain model without mixing. Only applying S2T mixing neglects the instances in target images. Through BDM, we achieve a balance between learning target features and mitigating the negative impact of incorrect target pseudo-labels on the model. We also perform hyper-parameters searching for the weights of the BDM loss when applying all three UDA methods. The results are shown in Table 4.2.1. The best result yields at the setting of = 0.4 and = 0.6, however, the results are stable for the different assessed combinations.
4.2.3 Category Module Training.
As mentioned, we perform CMT to introduce inductive bias, hoping to guide the model in better learning class features. To validate our conjecture, in the UrbanSyn Cityscapes setting, we perform CMT on two category groups, which are random compositions of classes (setting 2-5 in Table 5). Setting 1 is for full-class training. Overall, we observe that all four CMT models outperform the full-class training one. The best setting is 5, which achieves the highest improvement of +3.5 mAP compared to the model without CMT. We take this setting as the default for the rest of the experiments. However, three settings out of four show stable results. As mentioned in Sec. 3, we offer an alternative that enables the use of a single model during the testing phase. The category-free results for this approach are presented in Table 4.
| AP | AP50 | |||||||||||||||||||
| Model | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP | Diff. | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP50 | Diff. |
| UDA4Inst | 30.0 | 26.4 | 49.2 | 42.9 | 64.4 | 58.1 | 21.3 | 19.7 | 39.0 | - | 60.1 | 60.7 | 75.1 | 56.4 | 78.0 | 77.8 | 48.7 | 53.6 | 63.8 | - |
| category-free | 30.4 | 26.6 | 48.8 | 46.3 | 63.0 | 51.6 | 22.7 | 19.4 | 38.6 | -0.4 | 63.1 | 62.4 | 75.7 | 57.3 | 76.3 | 80.1 | 50.9 | 54.8 | 65.1 | +1.3 |
4.2.4 Component Ablation.
To assess the impact of UDA4Inst components—BDM, RcB, and CMT—we perform an ablation study in the UrbanSyn Cityscapes setting in Table 6. The ablative models are named from M1 to M5 (from top to bottom) for illustration. Overall, M1 without UDA obtains the lowest mAP of 32.0, while M6, equipped with all three UDA methods, achieves the highest score of 39.0. When comparing M1 and M2, applying CMT can effectively improve the model performance by +2.8 mAP. The conclusion is further strengthened through the comparison of M3 and M4. The effect of BDM is demonstrated by the performance jump from 32.0 to 36.1 mAP (+4.1) from M1 to M3. Moreover, under the same circumstances as CMT, BDM helps to improve the precision in M4 across most classes compared to M2, contributing to a +2.9 mAP improvement. The RcB method, which is based on BDM and CMT, is applied in M5. Focusing on the defined rare classes, RcB helps M5 to obtain +10.8 and +1.8 AP gains to the M4 in ‘Train’ and ‘Motorcycle’, respectively. M4, using BDM alone without RcB, scores -1.3 mAP lower than M5. Comparing ablative models, BDM stands out as the most beneficial method for domain adaptation.
| Setting | Category Groups | mAP | mAP50 |
| 1 | {Person, Rider, Motorcycle, Bicycle, Car, Truck, Bus, Train} | 32.0 | 54.4 |
| 2 | {Person, Rider, Bus, Motorcycle}, {Car, Truck, Train, Bicycle} | 33.9 | 57.8 |
| 3 | {Person, Rider, Car, Bicycle}, {Truck, Bus, Train, Motorcycle} | 35.0 | 59.4 |
| 4 | {Person, Rider}, {Car, Truck, Bus, Train, Motorcycle, Bicycle} | 35.1 | 59.3 |
| 5 | {Person, Rider, Motorcycle, Bicycle}, {Car, Truck, Bus, Train} | 35.5 | 59.4 |
| CMT | BDM | RcB | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP | |
| M1 | - | - | - | 25.1 | 19.5 | 39.8 | 37.6 | 51.5 | 46.7 | 19.2 | 16.2 | 32.0 |
| M2 | ✓ | - | - | 25.2 | 20.7 | 47.0 | 40.2 | 60.1 | 52.0 | 17.1 | 16.1 | 34.8 |
| M3 | - | ✓ | - | 27.7 | 23.2 | 44.2 | 42.3 | 61.7 | 52.7 | 21.2 | 15.5 | 36.1 |
| M4 | ✓ | ✓ | - | 28.9 | 27.6 | 50.0 | 43.6 | 63.7 | 47.3 | 19.5 | 20.7 | 37.7 |
| M5 | ✓ | ✓ | ✓ | 30.0 | 26.4 | 49.2 | 42.9 | 64.4 | 58.1 | 21.3 | 19.7 | 39.0 |
5 Conclusion
In this paper, we present a framework named UDA4Inst that aims to provide a strong baseline of synth-to-real UDA for instance segmentation. We incorporate three effective UDA methods into the vanilla Mask2Former training protocol: instance-level bidirectional data mixing, online rare-class balancing, and category module training. In a comprehensive evaluation of three synth-to-real domain settings: UrbanSyn/Synscapes/SYNTHIA Cityscapes, we show that UDA4Inst achieves significant improvements in instance segmentation compared to the source-only Mask2Former. This is the first work to provide competitive AP results on Synscapes and the recently published synthetic dataset UrbanSyn. Our proposed UDA4Inst model yields state-of-the-art results on SYNTHIA. We anticipate that this work will serve as a foundational framework for future research in UDA for instance segmentation.
5.0.1 Acknowledgements
This work has been supported by the Spanish grant Ref. PID2020-115734RB-C21 (ADA/SSL-ADA subproject) funded by MCIN / AEI / 10.13039 / 501100011033. Yachan Guo acknowledges the financial support to her PhD from the Chinese Scholarship Council (CSC), grant number 202208310071. Antonio M. Lopez acknowledges the financial support to his general research activities given by ICREA under the ICREA Academia Program. All authors acknowledge the support of the Generalitat de Catalunya CERCA Program and its ACCIO agency to CVC’s general activities.